# T-PRIME: Transformer-based Protocol Identification for Machine-learning at the Edge

### 09 Jan 2024 v1

@@ I Introduction
Legacy protocol detection methods are integrated into the RF receiver network interface card, enabling fast preamble correlations in hardware. However, updating the system for new protocols leads to backward compatibility issues and potential detection errors in challenging wireless channels. On the other hand, software-defined radio (SDR) systems offer flexibility but introduce higher latencies due to data transfer and complexities in I/O and buffer management. This paper demonstrates that using edge devices with CPU and GPU on a system-on-a-module (SOM), along with careful software design, can overcome the these limitations and enable real-time processing for complex ML wireless applications on an SDR-based edge device.
@@ II Related Work
$\bullet$ ML-based methods. Protocol classification via ML methods is still in a nascent stage. Zhang et al.  [1] perform protocol detection between WiFi, LTE, and 5G signals using a multi-layer bidirectional RNN. They use 512 IQ pairs with a sliding window that as inputs to their RNN. Similarly, Schmidt et al. use a CNN to classify between WiFi, Bluetooth, and ZigBee signals [2]. Jagannath and Jagannath [3] perform both modulation classification and signal classification via CNNs by distinguishing between a wide range of different protocols including AM Radio, Bluetooth, WiFi, and IEEE 802.15.4 (Zigbee). However, these approaches do not address the challenge of classifying similar protocols, such as different WiFi standards.
@@ IV-2 Training Procedure
Once the burst signals have been generated, we use a PyTorch data generator to train our models that operates on a per-training batch basis to dynamically load into memory each file individually and simulate transmission through a channel model at an SNR level randomly chosen among the following options for (a) channel models: TGn, TGax, Rayleigh, no channel impairment; and (b) a random SNR level defined between -30 dB and 30 dB, applied using Additive White Gaussian Noise (AWGN). We use a wrapper for Matlab’s channel models standard compliant implementation and employ Model-B multi-path delay profile for TGn and TGax with device distance of 3m, while for Rayleigh model we use a single Non-Line Of Sight (NLOS) path with an average path gain of -3 dB and delay of 1.5e-9 seconds.
@@
The proposed T-PRIME design was finalized using the same training procedure above and a hyperparameters tuning phase based on a grid search on the validation set. The chosen training configuration for transformers architectures is 5 training epochs, batch size of 122 and learning rate 2e-4. We found that a stack of 2 transformer layers is optimal, along with two distinct sets of values for sequence length and slice length that provided nearly perfect results. The first model, named Small (SM) Transformer, utilized an input sequence of $M=24$ and tokens with a length of $S=64$. The second, Large (LG) model, employed a sequence of $M=64$ tokens with a length of $S=128$ each. The resulting number of parameters of SM and LG architectures are $1.6$M and $6.8$M, respectively.
@@ IV-A Comparison with state-of-the-art legacy and ML methods
We first show how T-PRIME approach can effectively improve the state-of-the-art over legacy approaches. Specifically, we use the same static test set described in Sec. IV-1 for a simulation study that processes each test transmission using the correlation-based procedure specified in the standard [10]. We employ a set of multiple decoders provided by Matlab that use the legacy method to detect the following preambles formats: 1) VHT (i.e. 802.11n), 2) HE-SU (i.e. 802.11ax), 3) Non-HT (i.e. 802.11g/b). We then process our test transmissions through these functions for each channel model considered. For each test signal, we simulate 100 transmissions over random channel instances for each channel model and SNR levels considered in our evaluation procedure described before. To report the accuracy of legacy procedures, we consider the inverse Packet-Detection-Rate (i.e. 1 - PDR), defined as the number of correct preamble format detection over the total number of simulated packets transmissions, and compare it with accuracy of proposed architectures under the same statistical channel and SNR conditions. We evince from Fig. 3(a) that both our T-PRIME models drastically outperform the legacy detection mechanism, achieving near-perfect detection accuracy already at SNR $=0$ dB, while traditional methods completely fail to operate under such circumstances. This demonstrates how these models are still able to extract and learn powerful time-domain dependencies from received signals even in presence of strong noise or low reception power.
After establishing superiority of our approach over legacy methods, we compare our model to other state-of-the-art ML methods for wireless signal classification. While our problem formulation differs from modulation classification, it serves as a relevant example of architectures processing sequences of raw IQ samples for similar tasks. We compare T-PRIME against three state-of-the-art models in this domain: ResNet [15], AMCNet [18], and MCformer [8], using the hyperparameters reported as optimal by their respective authors; we also report optimal results after our own hyperparameter search in Appendix. For further comparison, we trained a 1D CNN model inspired by [12] with a wider input size and slightly larger set of parameters. We used a common training configuration for all models with a batch size of 512, $M=1$ for input generation, CrossEntropyLoss as the loss function, and a learning rate of 0.001 with exponential decay to 0.0001 when a plateau in the validation loss is reached. Table II lists all the models we compared, along with their input size and total number of parameters.
@@ VII Conclusion and future work
T-PRIME achieves near-perfect classification accuracy for 4 different WiFi protocols, outperforming legacy methods in challenging conditions. T-PRIME proves superior to existing state of the art ML models at protocol detection, especially with low SNR and overlapping signals. While our current architecture employs fixed-length inputs, Transformers excel with variable length inputs, as shown by successful NLP models like ChatGPT. In the future, we will adapt our ML block to accommodate variable length inputsAs part of our real-time implementation, we identified different bottlenecks and offered insights for solutions, which can benefit other AI/ML-based classification systems running on resource-constrained edge devices. Finally, we provide public access to our code and datasets at [9].

### 04 Mar 2024 v2

@@ I Introduction
Legacy protocol detection methods are integrated into the RF receiver network interface card, enabling fast preamble correlations in hardware. However, updating the system for new protocols leads to backward compatibility issues and potential detection errors in challenging wireless channels. On the other hand, software-defined radio (SDR) systems offer flexibility but introduce higher latencies due to data transfer and complexities in I/O and buffer management. This paper demonstrates that using edge devices with CPU and GPU on a system-on-a-module (SOM), along with careful software design, can overcome these limitations and enable real-time processing for complex ML wireless applications on an SDR-based edge device.
@@ II Related Work
$\bullet$ ML-based methods. Protocol classification via ML methods is still in a nascent stage. Zhang et al.  [1] perform protocol detection between WiFi, LTE, and 5G signals using a multi-layer bidirectional RNN. They use 512 IQ pairs with a sliding window as inputs to their RNN. Similarly, Schmidt et al. use a CNN to classify between WiFi, Bluetooth, and ZigBee signals [2]. Jagannath and Jagannath [3] perform both modulation classification and signal classification via CNNs by distinguishing between a wide range of different protocols including AM Radio, Bluetooth, WiFi, and IEEE 802.15.4 (Zigbee). However, these approaches do not address the challenge of classifying similar protocols, such as different WiFi standards.
@@ IV-2 Training Procedure
Once the burst signals have been generated, we use a PyTorch data generator to train our models that operates on a per-training batch basis to dynamically load into memory each file individually and simulate transmission through a channel model at an SNR level randomly chosen. We consider the following options for (a) channel models: TGn, TGax, Rayleigh, no channel impairment; and (b) a random SNR level defined between $-30$ dB and $30$ dB, applied using Additive White Gaussian Noise (AWGN). We develop a Python wrapper for Matlab’s channel models standard compliant implementation and employ Model-B multi-path delay profile for TGn and TGax with device distance of $3$m, while for Rayleigh model we use a single Non-Line Of Sight (NLOS) path with an average path gain of $-3$ dB and delay of $1.5$e-$9$ seconds.
@@
The proposed T-PRIME design was finalized using the same training procedure above and a hyperparameters tuning phase based on a grid search on the validation set. The chosen training configuration for transformers architectures is $5$ training epochs, batch size of $122$ and learning rate $2$e44-4- 4. We found that a stack of $2$ transformer layers is optimal, along with two distinct sets of values for sequence length and slice length that provided nearly perfect results. The first model, named Small (SM) Transformer, utilized an input sequence of $M=24$ and tokens with a length of $S=64$. The second, Large (LG) model, employed a sequence of $M=64$ tokens with a length of $S=128$ each. The resulting number of parameters of SM and LG architectures are $1.6$M and $6.8$M, respectively.
@@ IV-A Comparison with state-of-the-art legacy and ML methods
We first show how T-PRIME approach can effectively improve the state-of-the-art over legacy approaches. Specifically, we use the same static test set described in Sec. IV-1 for a simulation study that processes each test transmission using the correlation-based procedure specified in the standard [10]. We employ a set of multiple decoders provided by Matlab that use the legacy method to detect the following preambles formats: 1) VHT (i.e. 802.11n), 2) HE-SU (i.e. 802.11ax), 3) Non-HT (i.e. 802.11g/b). We then process our test transmissions through these functions for each channel model considered. For each test signal, we simulate $100$ transmissions over random channel instances for each channel model and SNR levels considered in our evaluation procedure described before. To report the accuracy of legacy procedures, we consider the inverse Packet-Detection-Rate (i.e. $1$ PDR), defined as the ratio of correct preamble format detections over the total number of simulated packets transmissions, and compare it with accuracy of proposed architectures under the same statistical channel and SNR conditions. We evince from Fig. 3(a) that both our T-PRIME models drastically outperform the legacy detection mechanism, achieving near-perfect detection accuracy already at SNR $=0$ dB, while traditional methods completely fail to operate under such circumstances. This demonstrates how these models are still able to extract and learn powerful time-domain dependencies from received signals even in presence of strong noise or low reception power.
After establishing superiority of our approach over legacy methods, we compare our model to other state-of-the-art ML methods for wireless signal classification. While our problem formulation differs from modulation classification, it serves as a relevant example of architectures processing sequences of raw IQ samples for similar tasks. We compare T-PRIME against three state-of-the-art models in this domain: ResNet [15], AMCNet [18], and MCformer [8], using the hyperparameters reported as optimal by their respective authors; we also present supplementary findings obtained through additional hyperparameter exploration in Appendix. For further comparison, we trained a 1D CNN model inspired by [12] with a wider input size and slightly larger set of parameters. We used a common training configuration for all models with a batch size of 512, $M=1$ for input generation, CrossEntropyLoss as the loss function, and a learning rate of $0.001$ with exponential decay to $0.0001$ when a plateau in the validation loss is reached. Table II lists all the models we compared, along with their input size and total number of parameters.
@@ VII Conclusion and future work
T-PRIME achieves near-perfect classification accuracy for $4$ different WiFi protocols, outperforming legacy methods in challenging conditions. T-PRIME proves superior to existing state of the art ML models at protocol detection, especially with low SNR and overlapping signals. While our current architecture employs fixed-length inputs, Transformers excel with variable length inputs, as shown by successful NLP models like ChatGPT. In the future, we will adapt our ML block to accommodate variable length inputs. As part of our real-time implementation, we identified different bottlenecks and offered insights for solutions, which can benefit other AI/ML-based classification systems running on resource-constrained edge devices. Finally, we provide public access to our code and datasets at [9].
@@ VIII-B Test Results on Channel Models
VIII-C Detailed accuracy metrics for multi-class classification on overlapping datasets
In this section, we present a detailed analysis of the classification accuracy of T-PRIME LG across various overlapping transmission configurations outlined in Table IV. We introduce three distinct metrics for these results, one of which is innovative compared to the metrics presented in Table VI.
We emphasize two specific accuracy metrics: Exact Accuracy and Single Accuracy. Exact Accuracy pertains to precisely identifying all and only the protocols being transmitted, while single accuracy denotes the percentage of transmissions where we successfully detect at least one of the transmitted protocols, regardless of any incorrect detection of protocols not present in the signal.
The new metric, referred to as Single Exact Accuracy, quantifies instances in which we detect at least one of the transmitted protocols while simultaneously avoiding the detection of any non-transmitted protocols. Specifically, this metric recognizes as correct only those samples containing either one or both of the transmitted protocols, counting a sample as incorrect if classifier outputs any other protocol combination.
TABLE XVII: Global results
TABLE XVIII: Dataset RM_C_O1 (overlap ratio $R=25\%$)
TABLE XIX: Dataset RM_C_O1 (overlap ratio $R=50\%$)
TABLE XX: Dataset RM_C_O2 (overlap ratio $R=25\%$)
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 69.9 63.8 69.4 62.5 76.9 79.5
Single Exact Accuracy 80.2 76.5 79.4 70.7 82.2 82.6
Single Accuracy 99.8 100.0 99.9 99.9 100.0 100.0
ax detected 80.8 6.4 9.2 92.9 90.6 16.5
b detected 92.9 100.0 96.4 7.0 2.7 0.8
n detected 13.8 17.3 77.7 75.1 15.2 84.5
g detected 6.2 65.7 11.6 22.3 88.9 97.1
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 72.6 97.9 90.9 38.9 84.8 89.1
Single Exact Accuracy 79.6 98.8 94.3 53.1 90.9 94.3
Single Accuracy 99.8 100.0 100.0 99.8 99.9 100.0
ax detected 92.4 0.2 1.5 97.8 99.8 4.0
b detected 84.7 100.0 95.4 19.7 7.6 1.7
n detected 16.5 1.0 96.2 45.9 1.5 98.2
g detected 4.0 98.0 4.2 27.2 86.2 92.7
noise detected 0.0 0.0 0.0 0.1 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 82.8 58.8 70.5 40.3 69.6 96.3
Single Exact Accuracy 87.5 64.0 75.6 51.8 76.3 97.3
Single Accuracy 99.7 100.0 99.9 100.0 100.0 99.9
ax detected 93.1 1.4 5.1 79.8 95.9 1.6
b detected 94.0 100.0 93.3 2.3 1.2 1.1
n detected 12.3 34.7 85.9 74.2 22.5 98.8
g detected 0.3 62.6 19.3 45.9 78.4 98.4
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 49.5 33.4 28.1 97.6 71.4 55.0
Single Exact Accuracy 70.7 58.3 57.1 98.6 76.8 58.7
Single Accuracy 100.0 100.0 99.8 99.9 100.0 100.0
ax detected 53.3 20.8 26.7 99.1 72.6 41.3
b detected 99.7 100.0 99.7 1.2 0.0 0.0
n detected 10.4 21.8 31.3 98.5 23.2 57.7
g detected 19.2 35.0 17.2 0.3 99.7 99.5
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
Overlapping configuration (incumbent - interferer)
Metrics (%) b - ax b - g b - n n - ax g - ax g - n
Exact Accuracy 64.6 49.4 75.6 97.6 81.5 63.2
Single Exact Accuracy 79.5 79.4 84.5 99.1 83.2 65.9
Single Accuracy 99.8 100.0 100.0 100.0 99.9 100.0
ax detected 71.4 9.3 10.1 99.5 86.4 34.1
b detected 96.9 100.0 99.5 0.4 0.0 0.0
n detected 15.2 11.3 83.0 98.2 16.8 68.3
g detected 5.6 51.4 5.3 0.4 98.7 99.7
noise detected 0.0 0.0 0.0 0.0 0.0 0.0
TABLE XVII: Global results
TABLE XVIII: Dataset RM_C_O1 (overlap ratio $R=25\%$)
TABLE XIX: Dataset RM_C_O1 (overlap ratio $R=50\%$)
TABLE XX: Dataset RM_C_O2 (overlap ratio $R=25\%$)
TABLE XXI: Dataset RM_C_O2 (overlap ratio $R=50\%$)
From Table XXI we observe an average of 86.88% correct detection rates for protocol classes that are actually present within the overlapping signals in the test set, while detection rate for protocols not present is on average 7.16%. This demonstrates how our classifier can accurately identify the protocols even in presence of active interference, and it is able to correctly determine both incumbent and interferer transmissions in the majority of cases.
Upon a more detailed analysis of the results, looking at Tables XXI, XXI, XXI and XXI it becomes evident that 802.11b is the protocol most easily identified by the model. This behavior may arise from a number of factors, including distinctiveness of its waveforms compared to others and oversampling applied to signals in this category in order to match other protocols’ configurations in the rest of the dataset.
Another noteworthy observation is the distinct behavior of the model across the two datasets. In dataset RM_C_02, we generally observe superior detection accuracy for the incumbent signal compared to the interferer. However, in dataset RM_C_01, the model deviates from this pattern, exhibiting numerous instances where higher accuracy is achieved for the interferer.
Finally, it is noteworthy to observe the model’s difficulty in identifying the 802.11g protocol particularly when mixed as an interferer with 802.11b. In this specific configuration, the former protocol presents the most significant challenge for the model in terms of identification. Nevertheless, the model overall continues to yield satisfactory results, motivating further research efforts in these overlapping scenarios.